OpenStudy (anonymous):

Prove: sqrt(cosec^A-1) = cosAcosecA

OpenStudy (anonymous):

\[\sqrt{cosec^2A-1}=cosAcosecA\]

hartnn (hartnn):

use, cosec^2 = 1+cot^2

OpenStudy (anonymous):

\[\sqrt{(1+\cot^2A)-1}\]

OpenStudy (anonymous):

\[\sqrt{\cot^2A}\]

OpenStudy (anonymous):

??

hartnn (hartnn):

yes. =cot = cos/sin and 1/sin =cosec

OpenStudy (anonymous):

use \[\frac{ 1 }{ \sin ^{2}A } = \csc ^{2}A\] to replace the cosec^2A

OpenStudy (anonymous):

After, you can find the LCM which is sin^2 A.

OpenStudy (anonymous):

Realize you get a numerator which is 1-sin^2 A which can be replaced by cos^2 A using the identity: \[\cos ^{2}A + \sin ^{2}A = 1\]

OpenStudy (anonymous):